OpenStudy (anonymous):

Show that the series Σ1/(1+xn^2) converges uniformly on [a,∞) for any a>0 but does not converge uniformly on (0,∞).

OpenStudy (blockcolder):

\[\sum 1+(1+xn^2)\] is the series?

OpenStudy (anonymous):

yes

OpenStudy (blockcolder):

But that doesn't converge for any value of x because \(\lim_{n \to \infty} (2+xn^2)=\infty\) for all x.

OpenStudy (anonymous):

Um, she said yes, but I think that's not the series. The series she typed in her first post is \[\sum \frac{1}{1+xn^2}\]

OpenStudy (blockcolder):

That's because she edited her post.

OpenStudy (anonymous):

lols